OpenStudy (anonymous):

Solve the logarithmic equation (show all work) (Do NOT use a calculator) 1. Log6x = 1 – log6(x-5) 2. 2x^2 – 9x = 1/256 3. 9x-1 = 100(3x) 4. 2lnx – ln5 = ln(x + 10) 5. ex – 2 = -e –x 6. 2log50 = 3log25 + log(x-2)

OpenStudy (anonymous):

Here are some of the rules for ln's (natural logarithms)(or logs in general). log(xy) = log x + log y log(x/y) = logx - log y log x^p = p log x